4. EX08-06.ALGO (Algorithmic) Prowiding for Doubtful Accounts At the end of the current yea, the accounts recelvable account has a debit balance of $755,000 and sales for the year total $8,560,000, a. The allowance account before adjustment has a debit balance of $10,200. Bad debt expense ls estimated at 1/4 of 1% of sales. b. The allowance account before adjustment has a deblt balance of $10,200. An aging of the accounts in the oustomer ledger indlcates estimated doubtful accounts of $32,600 c. The allowance account before adjustment has a credit balance of $8,200. Bad debt expense ls estimated at 3/4 of 1% of sales. d. The allawance account before adjustment has a credit balance of $8,200, An aging of the accounts in the oustomer ledger indicates estimated doubtful accounts of $68,100 Determine the amount of the adjusting entry to prcwide for doubtfuli accounts under each of the assumptions (a through d) isted above d.

8-2-4
Answer #1
a (8,560,000*1% * 1/4) - 10,200 = 21,400-10,200 = 11,200
b 32,600 + 10,200 = 42,800
c (8,560,000*1%*3/4) - 8,200 = (64,200-8,200) = 56,000
d (68,100-8,200) = 59,900